# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ D.3.29 Who Is Rich, Klam
Figure 69: Who Is Rich, Klam [72]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.30 After I’m Gone, Lippman
@@
Figure 71: After I’m Gone, Lippman [80]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.31 Sunburn, Lippman
Figure 73: Sunburn, Lippman [81]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.32 A Game of Thrones, Martin
@@ A Game of Thrones
Figure 75: A Game of Thrones, Martin [85]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.33 Twilight, Meyer
Figure 77: Twilight, Meyer [87]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.34 Beloved, Morrison
Figure 79: Beloved, Morrison [89]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.35 The Memory Police, Ogawa
@@
Figure 81: The Memory Police, Ogawa [96]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.36 Nineteen Eighty-Four, Orwell
@@
Figure 83: Nineteen Eighty-Four, Orwell [100]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.37 The Subtle Knife, Pullman
@@
Figure 85: The Subtle Knife, Pullman [102]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.38 The Making of a Mediterranean Emirate, Rouighi
@@
Figure 87: The Making of a Mediterranean Emirate, Rouighi [105]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.39 Harry Potter and the Sorcerer’s Stone, Rowling
@@
Figure 89: Harry Potter and the Sorcerer’s Stone, Rowling [106]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.40 Harry Potter and the Goblet of Fire, Rowling
@@
Figure 91: Harry Potter and the Goblet of Fire, Rowling [107]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.41 The Catcher in the Rye, Salinger
Figure 93: The Catcher in the Rye, Salinger [110]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.42 Lean In, Sandberg
Figure 95: Lean In, Sandberg [112]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.43 The Bedwetter, Silverman
Figure 97: The Bedwetter, Silverman [117]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.44 No Visible Bruises, Snyder
@@
Figure 99: No Visible Bruises, Snyder [118]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.45 Unglued, TerKeurst
Figure 101: Unglued, TerKeurst [122]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.46 Embraced, TerKeurst
Figure 103: Embraced, TerKeurst [123]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.47 The Hobbit, Tolkien
Figure 105: The Hobbit, Tolkien [126]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.48 Brown Girl Dreaming, Woodson
@@
Figure 107: Brown Girl Dreaming, Woodson [134]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.49 Another Brooklyn, Woodson
@@
Figure 109: Another Brooklyn, Woodson [135]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.50 The Future of the Internet and How to Stop It, Zittrain
Figure 111: The Future of the Internet and How to Stop It, Zittrain [139]: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.4.1 Quantifying extraction coverage
For details on $\tau_{\text{min}}$, see Equation 9 (Sections 2.15; Appendices CE). In the main text, we omit $\theta$ and $\phi$ for brevity, since we always specify the exact model and decoding scheme being considered. The results of our sliding-window procedure can be used to compute Equation 10. We pick a threshold $\tau$, then we filter the whole (sorted by start position) list of $p_{\bm{z}}$ we obtained with the sliding-window procedure to only contain sequences whose extraction probability is $\geq\!\tau$. Our sliding-window procedure keeps track of character-level metadata so that we can visualize heatmaps; we have information about the start/end character indexes for the prefixes and suffixes. Based on these intervals, we can compute which characters in the whole book are covered with $p_{\bm{z}}\!\geq\!\tau$, and then divide that by the total number of characters in the book. High extraction coverage does not necessarily mean we can extract the whole book in one go, or that we have identified the minimal set of sequences to try to extract to get maximal coverage of the book. This is simply a procedure for estimating the proportion of total memorized text in a book, in order to do cross-book and cross-model comparisons, as in Section 4.2.
@@ Appendix E Measurement validity
In this appendix, we provide more detail on the validity experiments introduced in Section 5. These experiments support that our measurements capture true instances of extraction. On this basis, we expand our analysis to models with undisclosed information about training on books—DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B. Because our controls show that our method does not mistake chance generation for extraction, we can be confident that when these models register extraction success, it reflects memorization. Given the length of this appendix, we provide a summary:
  • Extraction of known training data is evidence of memorization. For models where the training data are known, high-probability generation of a verbatim suffix given its prefix is widely accepted as evidence of memorization [16, 17, 18, 60]. Probabilistic extraction provides a more fine-grained version of the same test: rather than returning a single greedy continuation, it assigns a probability to the target suffix and reveals whether it surpasses a minimum acceptable threshold.
@@
  • Validating against non-training data The majority of our experiments—those on Llama and Pythia models—fit the case of extracting known training data from models. It is known that Books3 was included in the training data, and we are running experiments to extract Books3 sequences. But we can also do more to ensure validity: we can test our procedure on known non-training data, and show that suffix probabilities fall below our chosen $\tau_{\text{min}}$. We (and Hayes et al. [60]) do this for an additional sanity check. In particular, we perform over $200$ negative control experiments (Section 5.2 & Appendix E.4). Work on greedy-decoded discoverable extraction often does not do this; it assumes validity [18] or uses shorter prefix lengths to induce less bias on the output [93] (Appendix E.6). Overall, our results confirm that our choice of $\tau_{\text{min}}$ is very conservative; we could set it lower without risking false positives. However, we are deliberately cautious, and defer this to future work (Section 5.3).
  • Stress-testing extraction with longer prefixes. When validating $\tau_{\text{min}}$ on known training data, we can also ask whether longer prefixes make memorization more discoverable. Because suffix probabilities are conditional on the prefix, extending the prompt can bias the model’s output toward a certain part of its learned distribution (which can also further be affected by the decoding algorithm). As a result, as others have shown [18], for experiments on known training data, when a piece of training data is memorized, a longer prefix prompt can raise the probability of the target suffix and expose it as memorized (via greedy decoding). This memorization may not be discoverable with only a $50$-token prompt, but may become discoverable with a longer prompt. Our own experiments confirm these patterns, but also show clarify, for a specific work, that longer prompts do not always lead to successful extraction: if a sequence is not memorized, we are unable to find a long enough prefix to make it it extractable (Appendix E.3). This supports the conclusion that probabilistic extraction captures true memorization, and clarifies the limits of analogies that treat models as “monkeys at the typewriter” (Appendix E.5).

### v4

@@ D.3.29 Who Is Rich, rich
Figure 69: Who Is Rich, rich: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.30 After I’m Gone, afterimgone
@@
Figure 71: After I’m Gone, afterimgone: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.31 Sunburn, sunburn
Figure 73: Sunburn, sunburn: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.32 A Game of Thrones, thrones
@@ A Game of Thrones
Figure 75: A Game of Thrones, thrones: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.33 Twilight, twilight
Figure 77: Twilight, twilight: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.34 Beloved, beloved
Figure 79: Beloved, beloved: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.35 The Memory Police, memory
@@
Figure 81: The Memory Police, memory: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.36 Nineteen Eighty-Four, 1984
@@
Figure 83: Nineteen Eighty-Four, 1984: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.37 The Subtle Knife, subtleknife
@@
Figure 85: The Subtle Knife, subtleknife: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.38 The Making of a Mediterranean Emirate, emirate
@@
Figure 87: The Making of a Mediterranean Emirate, emirate: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.39 Harry Potter and the Sorcerer’s Stone, hp1
@@
Figure 89: Harry Potter and the Sorcerer’s Stone, hp1: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.40 Harry Potter and the Goblet of Fire, hp4
@@
Figure 91: Harry Potter and the Goblet of Fire, hp4: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.41 The Catcher in the Rye, catcher
Figure 93: The Catcher in the Rye, catcher: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.42 Lean In, leanin
Figure 95: Lean In, leanin: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.43 The Bedwetter, bedwetter
Figure 97: The Bedwetter, bedwetter: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.44 No Visible Bruises, novisible
@@
Figure 99: No Visible Bruises, novisible: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.45 Unglued, unglued
Figure 101: Unglued, unglued: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.46 Embraced, embraced
Figure 103: Embraced, embraced: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.47 The Hobbit, hobbit
Figure 105: The Hobbit, hobbit: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.48 Brown Girl Dreaming, browngirldreaming
@@
Figure 107: Brown Girl Dreaming, browngirldreaming: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@
D.3.49 Another Brooklyn, anotherbrooklyn
@@
Figure 109: Another Brooklyn, anotherbrooklyn: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.3.50 The Future of the Internet and How to Stop It, zittrain
Figure 111: The Future of the Internet and How to Stop It, zittrain: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
@@ D.4.1 Quantifying extraction coverage
For details on $\tau_{\text{min}}$, see Equation 9 (Sections 2.1 & 5; Appendices C & E). In the main text, we omit $\theta$ and $\phi$ for brevity, since we always specify the exact model and decoding scheme being considered.
The results of our sliding-window procedure can be used to compute Equation 10. We pick a threshold $\tau$, then we filter the whole (sorted by start position) list of $p_{\bm{z}}$ we obtained with the sliding-window procedure to only contain sequences whose extraction probability is $\geq\!\tau$. Our sliding-window procedure keeps track of character-level metadata so that we can visualize heatmaps; we have information about the start/end character indexes for the prefixes and suffixes. Based on these intervals, we can compute which characters in the whole book are covered with $p_{\bm{z}}\!\geq\!\tau$, and then divide that by the total number of characters in the book.
High extraction coverage does not necessarily mean we can extract the whole book in one go, or that we have identified the minimal set of sequences to try to extract to get maximal coverage of the book. This is simply a procedure for estimating the proportion of total memorized text in a book, in order to do cross-book and cross-model comparisons, as in Section 4.2.
@@ Appendix E Measurement validity
In this appendix, we provide more detail on the validity experiments introduced in Section 5. These experiments support that our measurements capture true instances of extraction. On this basis, we expand our analysis to models with undisclosed information about training on books—DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B. Because our controls show that our method does not mistake chance generation for extraction, we can be confident that when these models register extraction success, it reflects memorization.
Given the length of this appendix, we provide a summary:
  • Extraction of known training data is evidence of memorization. For models where the training data are known, high-probability generation of a verbatim suffix given its prefix is widely accepted as evidence of memorization [carlini2021extracting, carlini2023extracting, carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels]. Probabilistic extraction provides a more fine-grained version of the same test: rather than returning a single greedy continuation, it assigns a probability to the target suffix and reveals whether it surpasses a minimum acceptable threshold.
@@
  • Validating against non-training data The majority of our experiments—those on Llama and Pythia models—fit the case of extracting known training data from models. It is known that Books3 was included in the training data, and we are running experiments to extract Books3 sequences. But we can also do more to ensure validity: we can test our procedure on known non-training data, and show that suffix probabilities fall below our chosen $\tau_{\text{min}}$. We (and hayes2025measuringmemorizationlanguagemodels) do this for an additional sanity check. In particular, we perform over $200$ negative control experiments (Section 5.2 & Appendix E.4). Work on greedy-decoded discoverable extraction often does not do this; it assumes validity [carlini2023quantifying] or uses shorter prefix lengths to induce less bias on the output [nasr2023scalable] (Appendix E.6). Overall, our results confirm that our choice of $\tau_{\text{min}}$ is very conservative; we could set it lower without risking false positives. However, we are deliberately cautious, and defer this to future work (Section 5.3).
  • Stress-testing extraction with longer prefixes. When validating $\tau_{\text{min}}$ on known training data, we can also ask whether longer prefixes make memorization more discoverable. Because suffix probabilities are conditional on the prefix, extending the prompt can bias the model’s output toward a certain part of its learned distribution (which can also further be affected by the decoding algorithm). As a result, as others have shown [carlini2023quantifying], for experiments on known training data, when a piece of training data is memorized, a longer prefix prompt can raise the probability of the target suffix and expose it as memorized (via greedy decoding). This memorization may not be discoverable with only a $50$-token prompt, but may become discoverable with a longer prompt.
    Our own experiments confirm these patterns, but also show clarify, for a specific work, that longer prompts do not always lead to successful extraction: if a sequence is not memorized, we are unable to find a long enough prefix to make it it extractable (Appendix E.3). This supports the conclusion that probabilistic extraction captures true memorization, and clarifies the limits of analogies that treat models as “monkeys at the typewriter” (Appendix E.5).